A pharmacist adds pure alcohol to a 25% alcohol solution to obtain 6 quarts of a 45% solution of alcohol. How much of the 25% solution and how much pure alcohol did she use?
Answer by mananth(16946)

---------------- percent ---------------- quantity
Solution I 25% ---------------- x
Solution II 100% ---------------- 6 - x
Mixture 45.00% ---------------- 6
0.25x+1( 6 - x ) = 6.00 * 0.45

0.25x+6 -x = 2.70
0.25x -1x = 2.7 - 6
-0.75 x = -3.3

/ -0.75
x =4.4quarts 25.00% Alcohol
1.6 quarts 100.00% pure alcohol